# Spinning LQG black hole as a particle accelerator

### v1

@@ Abstract
Loop Quantum Gravity, BSW Mechanism, Black hole particle accelerator
@@ I Introduction
Since their inception, black holes have been among the most intriguing objects in physics. While the theoretical aspects of black holes have been extensively explored for a long time, the observational confirmation and validation of these theoretical predictions have only been achieved recently. The LIGO-Virgo-KAGRA collaboration’s detection of gravitational waves originating from the collision of compact binary objects provided the first direct evidence for the existence of black holes, opening a new field of research known as black hole spectroscopy (Abbott:2016blz, ; TheLIGOScientific:2016src, ; Abbott:2016nmj, ). Furthermore, the Event Horizon Telescope (EHT) succeeded in imaging the shadow of a black hole, an achievement made possible through the observation of gravitational light bending and photon capture at the event horizons of supermassive black holes in the center of the giant elliptical galaxy M87 and at the center of our galaxy, Sgr A (Akiyama:2019bqs, ; Akiyama:2019cqa, ; Akiyama:2019fyp, ; EventHorizonTelescope:2022wkp, ; EventHorizonTelescope:2022xqj, ). Although these observations align with the predictions of general relativity with remarkable precision, fundamental questions remain unresolved. One such question pertains to the validity of the cosmic censorship conjecture, which says that spacetime singularities are always hidden behind an event horizon, thereby preventing the existence of naked singularities. In the context of black hole physics, these questions are of significant importance, as black holes are strongly gravitating objects that may possess spacetime singularities at their centers. However, the combination of gravity with quantum theory could potentially mitigate the occurrence of curvature singularities in the Universe.
@@
Spinning black hole solutions are of significant interest because most astrophysical black holes are characterized by definite angular momentum EventHorizonTelescope:2022xqj . However, it should be noted that studies of black hole singularities in LQG are mostly confined to spherically symmetric spacetimes due to their relative simplicity. Nevertheless, a spinning LQG black hole solution was constructed in Ref. Kumar:2022vfg using a modified Newman-Janis algorithm. In addition to the conventional black hole parameters, such as mass and angular momentum, the spinning LQG Kerr-like solution introduces an additional parameter, $\ell$, arising from quantum effects. It was shown that this additional parameter can be constrained by studying the shadow cast by M87 and Sgr A as observed by the Event Horizon Telescope Afrin:2022ztr ; Islam:2022wck ; Kumar:2023jgh .
@@
A notable feature of spinning black holes is the possibility of extracting mechanical energy from them. The concept of extracting the rotational energy of a black hole, originally introduced by Penrose, is known as the Penrose process Penrose:1971uk . In this process, under certain physical conditions, a particle scattered by a black hole can emerge with more energy than it originally had, with the extra energy derived from the spin of the black hole. This process is particularly efficient in scenarios where two particles collide near the event horizon of the black hole, producing two or more resultant particles. One of these particles may escape, carrying more energy than the incoming ones; this phenomenon is referred to as the collisional Penrose process Piran1977 . Interest in this concept was rejuvenated with the introduction of the Bañados-Silk-West (BSW) mechanism Banados:2009pr . This mechanism suggests that rotating black holes can function as natural particle accelerators, potentially explaining highly energetic astrophysical phenomena such as active galactic nuclei, gamma-ray bursts, and ultra-high-energy cosmic rays. The BSW mechanism significantly increases the efficiency of energy extraction, potentially reaching energy levels on the order of Planck-scale physics Bejger:2012yb ; Schnittman:2014zsa ; Berti:2014lva ; Leiderschneider:2015kwa . In the BSW mechanism, the center-of-mass energy of colliding particles can become exceedingly high near the event horizon of an extremal black hole, particularly when one particle has critical angular momentum. Due to its theoretical and observational importance, this mechanism has been the subject of extensive research Berti:2009bk ; Banados:2010kn ; Jacobson:2009zg ; Zaslavskii:2010aw ; Wei:2010gq ; Lake:2010bq ; Wei:2010vca ; Liu:2010ja ; Mao:2010di ; Zhu:2011ae ; Zaslavskii:2012fh ; Zaslavskii:2012qy ; Zaslavskii:2010pw ; Grib:2010xj ; Harada:2011xz ; Liu:2011wv ; Patil:2010nt ; Patil:2011aw ; Patil:2011ya ; Patil:2011uf ; Amir:2015pja ; Ghosh:2014mea ; NaveenaKumara:2020kpz ; AhmedRizwan:2020sza ; Nosirov:2024vsu (see Ref. Harada:2014vka for a brief review).
@@ II Spinning LQG black hole
$ds^{2}=-\left(1-\frac{2M(r)\sqrt{r^{2}+\ell^{2}}}{\rho^{2}}\right)dt^{2}+\frac{\rho^{2}}{\Delta}dr^{2}+\rho^{2}d\theta^{2}-\frac{4aM(r)\sqrt{r^{2}+\ell^{2}}\sin^{2}\theta}{\rho^{2}}dtd\varphi+\frac{\mathcal{A}\sin^{2}\theta}{\rho^{2}}d\varphi^{2}$ (1)
@@
Here, $M$ is the black hole mass, $a$ is the black hole spin parameter, and $\ell$ is the polymer length scale, a positive constant with a dimension of length. The metric is asymptotically flat as $r\to\infty$. Additionally, the black hole solution exists for all values of $\ell$. Although the metric resembles the well-known Kerr black hole solution, it captures important features of LQG, such as a transition surface at the black hole center and the global regularity of spacetime. We present a brief overview of the spacetime properties, focusing mainly on the horizon structure and ergoregion. For further details on the spacetime structure, including the Penrose diagram, see Ref. Kumar:2022vfg . Despite its complex spacetime structure, the metric possesses notable limiting cases. In the limit $\ell\to 0$, we recover the Kerr black hole spacetime; in the limit $a\to 0$, we obtain spherical LQG black holes; and when both $a$ and $\ell$ vanish, we recover the Schwarzschild spacetime. Thus, the polymer length scale $\ell$ is key to the rich spacetime structure.
@@
The metric has divergences at $\Delta=0$, indicating a coordinate singularity corresponding to the black hole horizon. The location of this null surface is determined by the LQG parameter $\ell$ as well as the conventional parameters $a$ and $M$. The classical ring singularity of the Kerr black hole is replaced by a timelike transition surface at $r=0$ inside the event horizon. Another contrasting feature to the Kerr spacetime is the multihorizon structure. The parameter space of $\ell$ and $a$ produces three distinct scenarios (excluding naked singularities): black holes with one ($r_{1}$), two ($r_{1}>r_{2}$), or three horizons ($r_{1}>r_{2}>r_{3}$), where these are the positive roots of the condition $\Delta=0$. Here, $r_{1}\equiv r_{\text{h}}$ is the event horizon, $r_{2}\equiv r_{\text{c}}$, if it exists, is the Cauchy horizon, and $r_{3}$ (conditionally present) is an additional horizon inside the Cauchy horizon. For fixed values of $a$ or $\ell$, there exists a critical value of $\ell$ or $a$ for which the Cauchy horizon and event horizon merge, resulting in a maximally spinning black hole. This merging is characterized by the simultaneous solution of the equations $\Delta=0$ and $\Delta^{\prime}=0$
@@
For a given $\ell$, we can calculate the radius of the event horizon $r_{\text{he}}$ and the spin $a_{\text{e}}$ of the extremal black hole
@@
Although black hole solutions may exist for arbitrary values of the parameter $\ell$, extremal solutions are guaranteed only within a specific range of $\ell$ values. In the range $0\leq\ell\leq 0.60057$, the scenario is analogous to Kerr spacetime, where one finds non-extremal black holes, extremal black holes, and naked singularity solutions depending on the value of $a$. Specifically, for a given value of $\ell$, a non-extremal black hole becomes extremal when $a=a_{\text{e}}$, and for $a>a_{\text{e}}$, the solutions correspond to naked singularities. The extremal solution at $\ell=0.60057$ represents an extremal black hole with a null throat, characterized by $(a_{\text{e}}=1.09596,\ell=0.60057)$. In the domain $0.60057<\ell\leq 0.70711$, black hole solutions persist even for values of $a$ slightly greater than $a_{\text{e}}$, within a narrow region, before they transition into naked singularities. The extremal black hole solution at $\ell_{\text{e}}=0.70711$, called ultraextremal black hole, is characterized by $(a_{\text{e}}=1.13975,\ell_{\text{e}}=0.70711)$. For $\ell>0.70711$, extremal black hole solutions cease to exist (see Ref. Kumar:2022vfg for more details on this). Since the center-of-mass energy of colliding particles in the BSW mechanism increases indefinitely only in extremal cases, we restrict our consideration of the LQG parameter $\ell$ to the domain $0\leq\ell\leq 0.70711$.
@@ III Particle Motion around a Spinning LQG Black Hole
$\left(\frac{\partial S}{\partial\theta}\right)^{2}+\left(aE\sin{\theta}-\frac{L}{\sin{\theta}}\right)^{2}+m^{2}a^{2}\cos^{2}{\theta}=-\Delta\left(\frac{\partial S}{\partial r}\right)^{2}+\frac{[(a^{2}+\ell^{2}+r^{2})E-aL]^{2}}{\Delta}-m^{2}(r^{2}+\ell^{2}).$ (13)
@@ IV Particle Collisions near the Black Hole Horizon
$\frac{E^{2}_{CM}}{2m^{2}}=\left(1-g_{\mu\nu}\dot{x}_{1}^{\mu}\dot{x}_{2}^{\nu}\right).$ (34)
@@
$\Xi\equiv\frac{E_{\text{CM}}^{2}}{2m^{2}}=\frac{\left(a^{2}-aL_{1}+\ell^{2}+r^{2}\right)\left(a^{2}-aL_{2}+\ell^{2}+r^{2}\right)+\Delta\left(\ell^{2}+r^{2}-(a-L_{1})(a-L_{2})\right)-X_{1}X_{2}}{\Delta\left(\ell^{2}+r^{2}\right)}$ (36)
@@
The above equation remains unchanged if $L_{1}$ and $L_{2}$ are interchanged. This implies that the center-of-mass energy $E_{\text{CM}}$ is invariant under the exchange of particles (since we consider identical particles). We also note that the value of $E_{\text{CM}}$ deviates from the Kerr black hole case as it depends on the LQG parameter $\ell$, in addition to the black hole spin. This implies that the particle acceleration in the LQG black hole background is influenced by $\ell$. In Fig. 4, 5 and 6, we study the behavior of $E_{\text{CM}}$ as a function of the radial coordinate $r$ in the vicinity of the event horizon $r_{\text{h}}$, for extremal and non-extremal cases. The angular momenta of the particles are taken within the allowed range $L_{\text{min}}$ and $L_{\text{max}}$ so that they always approach the event horizon. Only particles possessing critical angular momentum can reach the horizon. However, it is important to note that the critical angular momentum, determined by Eq. (30), falls outside the range $(L_{\text{min}},L_{\text{max}})$ for non-extremal black holes, but it is within this range for extremal black holes. Without loss of generality, we set $m=1$ in our analysis.
For extremal black holes, as shown in Fig. 4 and 5, it is clear that the center-of-mass energy $E_{\text{CM}}$ remains finite when the angular momenta of the incoming particles are below the critical value, and it diverges when one of the particles has critical angular momentum. Although this result is similar to that of the Kerr black hole, the extremal spin in the presence of the LQG parameter can exceed the Kerr case, enhancing the potential to probe Planck-scale physics through particle acceleration. However, it is important to note that the extremal case is an idealization, as it takes an infinite amount of time for a particle to reach the horizon. Therefore, we expect that near-extremal cases can act as realistic particle accelerators, even though the time required is very long. We investigate the behavior of $E_{\text{CM}}$ for non-extremal black holes in Fig. 6, where it attains finite and sufficiently high values at the horizon in all cases.
@@
Figure 4: (a) Center-of-mass energy for the ultraextremal black hole ($a_{\text{e}}=1.13975$ and $\ell_{\text{e}}=0.70711$). (b) Center-of-mass energy for an extremal black hole with a null throat ($a_{\text{e}}=1.09596$ and $\ell=0.60057$).
@@
Figure 5: Center-of-mass energy for the extremal black hole. (a) $\ell=0.2$ and $a_{\text{e}}=1.01005$. (b) $\ell=0.4$ and $a_{\text{e}}=1.04091$
@@
Figure 6: Center-of-mass energy for the non-extremal black hole. (a) $a=1$ and $\ell=0.2$. (b) $a=1$ and $\ell=0.4$.

### v2

@@ Abstract
Loop Quantum Gravity, BSW Mechanism, Black hole particle accelerator.
@@ I Introduction
Since their inception, black holes have been among the most intriguing objects in physics. While the theoretical aspects of black holes have been extensively explored for a long time, the observational confirmation and validation of these theoretical predictions have only been achieved recently. The LIGO-Virgo-KAGRA collaboration’s detection of gravitational waves originating from the collision of compact binary objects provided the first direct evidence for the existence of black holes, opening a new field of research known as black hole spectroscopy Abbott:2016blz ; TheLIGOScientific:2016src ; Abbott:2016nmj . Furthermore, the Event Horizon Telescope (EHT) succeeded in imaging the shadow of a black hole, an achievement made possible through the observation of gravitational light bending and photon capture at the event horizons of supermassive black holes in the center of the giant elliptical galaxy M87 and at the center of our galaxy, Sgr A Akiyama:2019bqs ; Akiyama:2019cqa ; Akiyama:2019fyp ; EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj  [^1] A novel method is the Gaussian bending which could serve as a tool for testing various gravity theories and distinguishing physical objects with singularities from those without Zhang:2024uex ; Zhang:2021ygh . Although these observations align with the predictions of general relativity with remarkable precision, fundamental questions remain unresolved. One such question pertains to the validity of the cosmic censorship conjecture, which says that spacetime singularities are always hidden behind an event horizon, thereby preventing the existence of naked singularities. In the context of black hole physics, these questions are of significant importance, as black holes are strongly gravitating objects that may possess spacetime singularities at their centers. However, the combination of gravity with quantum theory could potentially mitigate the occurrence of curvature singularities in the Universe.
@@
Spinning black hole solutions are of significant interest because most astrophysical black holes are characterized by definite angular momentum EventHorizonTelescope:2022xqj . However, it should be noted that studies of black hole singularities in LQG are mostly confined to spherically symmetric spacetimes due to their relative simplicity. Nevertheless, a spinning LQG black hole solution was constructed in Ref. Kumar:2022vfg using a modified Newman-Janis algorithm. In addition to the conventional black hole parameters, such as mass and angular momentum, the spinning LQG Kerr-like solution introduces an additional parameter, $\ell$, arising from quantum effects. It was shown that this additional parameter can be constrained by studying the shadow cast by M87 and Sgr A as observed by the Event Horizon Telescope Afrin:2022ztr ; Islam:2022wck ; Kumar:2023jgh ; Vagnozzi:2022moj .
@@
A notable feature of spinning black holes is the possibility of extracting mechanical energy from them. The concept of extracting the rotational energy of a black hole, originally introduced by Penrose, is known as the Penrose process Penrose:1971uk . In this process, under certain physical conditions, a particle scattered by a black hole can emerge with more energy than it originally had, with the extra energy derived from the spin of the black hole. This process is particularly efficient in scenarios where two particles collide near the event horizon of the black hole, producing two or more resultant particles. One of these particles may escape, carrying more energy than the incoming ones; this phenomenon is referred to as the collisional Penrose process Piran1977 . Interest in this concept was rejuvenated with the introduction of the Bañados-Silk-West (BSW) mechanism Banados:2009pr . This mechanism suggests that rotating black holes can function as natural particle accelerators, potentially explaining highly energetic astrophysical phenomena such as active galactic nuclei, gamma-ray bursts, and ultra-high-energy cosmic rays. The BSW mechanism significantly increases the efficiency of energy extraction, potentially reaching energy levels on the order of Planck-scale physics Bejger:2012yb ; Schnittman:2014zsa ; Berti:2014lva ; Leiderschneider:2015kwa . In the BSW mechanism, the center-of-mass energy of colliding particles can become exceedingly high near the event horizon of an extremal black hole, particularly when one particle has critical angular momentum. Due to its theoretical and observational importance, this mechanism has been the subject of extensive research Berti:2009bk ; Banados:2010kn ; Jacobson:2009zg ; Zaslavskii:2010aw ; Wei:2010gq ; Lake:2010bq ; Wei:2010vca ; Liu:2010ja ; Mao:2010di ; Zhu:2011ae ; Zaslavskii:2012fh ; Zaslavskii:2012qy ; Zaslavskii:2010pw ; Grib:2010xj ; Harada:2011xz ; Liu:2011wv ; Patil:2010nt ; Patil:2011aw ; Patil:2011ya ; Patil:2011uf ; Amir:2015pja ; Ghosh:2014mea ; NaveenaKumara:2020kpz ; AhmedRizwan:2020sza ; Nosirov:2024vsu ; Galajinsky:2013as ; Tsukamoto:2013dna ; Ogasawara:2018gni (see Ref. Harada:2014vka for a brief review).
@@ II Spinning LQG black hole
$ds^{2}=$ $-\left(1-\frac{2M(r)\sqrt{r^{2}+\ell^{2}}}{\rho^{2}}\right)dt^{2}+\frac{\rho^{2}}{\Delta}dr^{2}+\rho^{2}d\theta^{2}-\frac{4aM(r)\sqrt{r^{2}+\ell^{2}}\sin^{2}\theta}{\rho^{2}}dtd\varphi$ (1)
$+\frac{\mathcal{A}\sin^{2}\theta}{\rho^{2}}d\varphi^{2}$
@@
Here, $M$ is the black hole mass, which we set to $M=1$ for simplicity wherever necessary. The spin parameter $a$ is related to the angular momentum $J$ and mass $M$ of the black hole as $a=J/M$, and $\ell$ is the polymer length scale, a positive constant  [^2] The polymer length scale represents a fundamental minimal length in polymer quantization, analogous to the discrete structures in loop quantum gravity Ashtekar:2002sn . This scale arises naturally from quantum geometry, highlighting the inherent discreteness of space, and plays a critical role in modifying the Hamiltonian of a system into a polymeric Hamiltonian.. In natural units, both $a$ and $\ell$ have the dimensions of length. The metric is asymptotically flat as $r\to\infty$. Additionally, the black hole solution exists for all values of $\ell$. Although the metric resembles the well-known Kerr black hole solution, it captures important features of LQG, such as a transition surface at the black hole center and the global regularity of spacetime. We present a brief overview of the spacetime properties, focusing mainly on the horizon structure and ergoregion. For further details on the spacetime structure, including the Penrose diagram, see Ref. Kumar:2022vfg . Despite its complex spacetime structure, the metric possesses notable limiting cases. In the limit $\ell\to 0$, we recover the Kerr black hole spacetime; in the limit $a\to 0$, we obtain spherical LQG black holes; and when both $a$ and $\ell$ vanish, we recover the Schwarzschild spacetime. Thus, the polymer length scale $\ell$ is key to the rich spacetime structure.
Figure 1: The plots depicting the different horizon structures of spinning LQG black holes, given by the roots of $\Delta=0$. Left: For a fixed $a$, varying $\ell$ changes $\Delta$ so that one, two, or three positive roots (i.e., horizons) may form, depending on whether $\Delta$ develops both a local maximum and minimum ($\ell<M/\sqrt{2}$) or remains strictly increasing ($\ell\geq M/\sqrt{2}$). Right: For a fixed $\ell$, increasing $a$ similarly alters the horizon structure.
@@
The metric has divergences at $\Delta=0$, indicating a coordinate singularity corresponding to the black hole horizon. The location of this null surface is determined by the LQG parameter $\ell$ as well as the conventional parameters $a$ and $M$. The classical ring singularity of the Kerr black hole is replaced by a timelike transition surface at $r=0$ inside the event horizon. Another contrasting feature to the Kerr spacetime is the multihorizon structure, as illustrated in Fig. 1. The function $\Delta$ governs the horizon structure and can exhibit up to three positive roots when $\ell<M/\sqrt{2}$, owing to a local maximum (at $r=(M-\sqrt{M^{2}-2\ell^{2}})/2$) and minimum (at $r=(M+\sqrt{M^{2}-2\ell^{2}})/2$). In this case, black holes may have multiple horizons. Conversely, for $\ell\geq M/\sqrt{2}$, $\Delta(r)$ is strictly increasing, admitting at most one horizon. The parameter space of $\ell$ and $a$ produces three distinct scenarios (excluding no-horizon spacetime): black holes with one ($r_{1}$), two ($r_{1}>r_{2}$), or three horizons ($r_{1}>r_{2}>r_{3}$), where these are the positive roots of the condition $\Delta=0$. Here, $r_{1}\equiv r_{\text{h}}$ is the event horizon, $r_{2}\equiv r_{\text{c}}$, if it exists, is the Cauchy horizon, and $r_{3}$ (conditionally present) is an additional horizon inside the Cauchy horizon. For fixed values of $a$ or $\ell$, there exists a critical value of $\ell$ or $a$ for which the Cauchy horizon and event horizon merge, resulting in a maximally spinning black hole. This merging is characterized by the simultaneous solution of the equations $\Delta=0$ and $\Delta^{\prime}=0$
@@
For a given $\ell$, we can calculate the extremal black hole’s radius of the event horizon $r_{\text{he}}$ and the extremal spin $a_{\text{e}}$ as
@@
Figure 2: The parameter space of $(a,\ell)$ for the spinning LQG black hole, illustrating regions of distinct horizon structures. BH I: single-horizon black holes, BH II: black holes with two horizons, BH III: black holes with three horizons, EBH: extremal black holes, NH: no-horizon spacetimes. The red curve indicates the critical $(a,\ell)$ values for which the event and Cauchy horizons coincide (extremal case), and beyond which there are no black hole solutions. The green line separates the BH I and BH III regions, while the blue line separates BH III from BH II and BH I from the NH regions.
The parameter-space diagram in Fig. 2 further demonstrates how varying $(a,\ell)$ leads to distinct black hole (BH) solutions, ranging from single- or multi-horizon configurations (BH I, BH II, BH III) to no-horizon (NH) spacetimes, and includes the orange line of extremal black holes (EBH) where horizons merge. Although black hole solutions may exist for arbitrary values of the parameter $\ell$, extremal solutions are guaranteed only within a specific range of $\ell$ values (orange line). In the range $0\leq\ell\leq 0.60057$, the scenario is analogous to Kerr spacetime, where one finds non-extremal black holes, extremal black holes, and no-horizon spacetimes depending on the value of $a$. Specifically, for a given value of $\ell$, a non-extremal black hole becomes extremal when $a=a_{\text{e}}$, and for $a>a_{\text{e}}$, the solutions correspond to no-horizon spacetimes. The extremal solution at $\ell=0.60057$ represents an extremal black hole with a null throat (green dot in Fig. 2), characterized by $(a_{\text{e}}=1.09596,\ell=0.60057)$. In the domain $0.60057<\ell\leq 0.70711$, black hole solutions persist even for values of $a$ slightly greater than $a_{\text{e}}$, within a narrow region, before they transition into no-horizon spacetimes. The extremal black hole solution at $\ell_{\text{e}}=0.70711$, called ultraextremal black hole (red dot in Fig. 2), is characterized by $(a_{\text{e}}=1.13975,\ell_{\text{e}}=0.70711)$. For $\ell>0.70711$, extremal black hole solutions cease to exist (see Ref. Kumar:2022vfg for more details on this). Since the center-of-mass energy of colliding particles in the BSW mechanism increases indefinitely only in extremal cases, we restrict our consideration of the LQG parameter $\ell$ to the domain $0\leq\ell\leq 0.70711$.
@@ III Particle Motion around a Spinning LQG Black Hole
$\left(\frac{\partial S}{\partial\theta}\right)^{2}+\left(aE\sin{\theta}-\frac{L}{\sin{\theta}}\right)^{2}+m^{2}a^{2}\cos^{2}{\theta}=$ $-\Delta\left(\frac{\partial S}{\partial r}\right)^{2}+\frac{[(a^{2}+\ell^{2}+r^{2})E-aL]^{2}}{\Delta}$ (13)
$-m^{2}(r^{2}+\ell^{2}).$
@@ IV Particle Collisions near the Black Hole Horizon
$\frac{E_{\text{CM}}^{2}}{2m^{2}}=\left(1-g_{\mu\nu}\dot{x}_{1}^{\mu}\dot{x}_{2}^{\nu}\right).$ (34)
@@
$\Xi$ $\equiv\frac{E_{\text{CM}}^{2}}{2m^{2}}$ (36)
$=\frac{\left(a^{2}-aL_{1}+\ell^{2}+r^{2}\right)\left(a^{2}-aL_{2}+\ell^{2}+r^{2}\right)+\Delta\left(\ell^{2}+r^{2}-(a-L_{1})(a-L_{2})\right)-X_{1}X_{2}}{\Delta\left(\ell^{2}+r^{2}\right)}$
@@
The above equation remains unchanged if $L_{1}$ and $L_{2}$ are interchanged. This implies that the center-of-mass energy $E_{\text{CM}}$ is invariant under the exchange of particles (since we consider identical particles). We also note that the value of $E_{\text{CM}}$ deviates from the Kerr black hole case as it depends on the LQG parameter $\ell$, in addition to the black hole spin. This implies that the particle acceleration in the LQG black hole background is influenced by $\ell$. We examine the characteristics of the center-of-mass energy as the radial coordinate $r$ nears the event horizon $r_{\text{h}}$ of the black hole. By taking the limit as $r\to r_{\text{h}}$ in the extremal scenario (applying L’Hôpital’s rule twice), we find
$\frac{E_{\text{CM}}^{2}}{2m^{2}}$ $=9.0151-A(L_{1},L_{2})-\frac{0.3043\cdot B(L_{1})\cdot B(L_{2})}{C(L_{1})\cdot C(L_{2})}+\frac{C(L_{2})\cdot D(L_{1})}{C(L_{1})\cdot F(L_{1})}+\frac{C(L_{1})\cdot D(L_{2})}{C(L_{2})\cdot F(L_{2})},$ (38)
where
$A(L_{1},L_{2})$ $=1.2672\,L_{1}+L_{1}\left(1.0490-1.0077\,L_{2}\right)-0.2182\,L_{2},$
$B(L_{z})$ $=7.5751-3.7985\,L_{z},$
$C(L_{z})$ $=\sqrt{4.3090+L_{z}\left(-4.3215+1.0835\,L_{z}\right)},$
$D(L_{z})$ $=-5.8901+L_{z}\left(6.7749+L_{z}\left(-0.3475+\left(-1.7915+0.5039\,L_{z}\right)L_{z}\right)\right),$
$F(L_{z})$ $=3.9769+L_{z}\left(-3.9884+\,L_{z}\right).$
Here, we have taken $\ell=0.4$ and $M=1$, along with the corresponding values for the angular momentum and horizon radius of the extremal black hole. Expression (38) diverges when one of the particles approaches the black hole with critical angular momentum $L_{\text{c}}$, as $C(L_{z})$ becomes zero in such cases. An analytical expression for the limiting values of the center-of-mass energy can be derived for any other value of $\ell$. However, it should be noted that this divergence occurs only for extremal black holes.
In Fig. 6, 7 and 8, we study the behavior of $E_{\text{CM}}$ as a function of the radial coordinate $r$ in the vicinity of the event horizon $r_{\text{h}}$, for extremal and non-extremal cases. The angular momenta of the particles are taken within the allowed range $L_{\text{min}}$ and $L_{\text{max}}$ so that they always approach the event horizon. Only particles possessing critical angular momentum can reach the horizon. However, it is important to note that the critical angular momentum, determined by Eq. (30), falls outside the range $(L_{\text{min}},L_{\text{max}})$ for non-extremal black holes, but it is within this range for extremal black holes. Without loss of generality, we set $m=1$ in our analysis.
@@
Figure 6: (a) Center-of-mass energy for the ultraextremal black hole ($a_{\text{e}}=1.13975$, $\ell_{\text{e}}=0.70711$, $r_{\text{he}}=0.5$, $L_{\text{c}}=1.79778$ and $L_{2}=L_{\text{min}}=-5.00926$). (b) Center-of-mass energy for an extremal black hole with a null throat ($a_{\text{e}}=1.09596$, $\ell=0.60057$, $r_{\text{he}}=0.76393$, $L_{\text{c}}=1.95755$ and $L_{2}=L_{\text{min}}=-4.95681$).
@@
Figure 7: Center-of-mass energy for the extremal black hole. (a) $\ell=0.2$, $a_{\text{e}}=1.01005$, $r_{\text{he}}=0.97958$, $L_{\text{c}}=1.99968$ and $L_{2}=L_{\text{min}}=-4.84256$. (b) $\ell=0.4$, $a_{\text{e}}=1.04091$, $r_{\text{he}}=0.91231$, $L_{\text{c}}=1.99421$ and $L_{2}=L_{\text{min}}=-4.88501$.
@@
Figure 8: Center-of-mass energy for the non-extremal black hole. (a) $a=1$, $\ell=0.2$, $r_{\text{h}}=1.12432$, $L_{\text{max}}=2.20028$ and $L_{2}=L_{\text{min}}=-4.83548$. (b) $a=1$, $\ell=0.4$, $r_{\text{h}}=1.22264$, $L_{\text{max}}=2.40120$ and $L_{2}=L_{\text{min}}=-4.85651$.
For extremal black holes, as shown in Fig. 6 and 7, it is clear that the center-of-mass energy $E_{\text{CM}}$ remains finite when the angular momenta of the incoming particles are below the critical value, and it diverges when one of the particles has critical angular momentum. Although this result is similar to that of the Kerr black hole, the extremal spin in the presence of the LQG parameter can exceed the Kerr case, enhancing the potential to probe Planck-scale physics through particle acceleration. However, it is important to note that the extremal case is an idealization, as it takes an infinite amount of time for a particle to reach the horizon. Therefore, we expect that near-extremal cases can act as realistic particle accelerators, even though the time required is very long.
We investigate the behavior of $E_{\text{CM}}$ for non-extremal black holes in Fig. 8, where it attains finite and sufficiently high values at the horizon in all cases. This behavior is explored through a numerical approach, where we calculate the center-of-mass energy for a near-extremal black hole with angular momentum $a=a_{\text{e}}-\epsilon$, where $\epsilon$ is a small parameter characterizing a slight deviation from the extremal case $a=a_{\text{e}}$. In our numerical computations, we use $L_{1}=L_{\text{min}}$ and $L_{2}=L_{\text{max}}$ in the expression for the center-of-mass energy (36) and take the limit $r\to r_{\text{h}}$, where $L_{\text{min}}$, $L_{\text{max}}$, and $r_{\text{h}}$ are numerically determined for $a=a_{\text{e}}-\epsilon$. The results, presented in Table 3, consider smaller values of $\epsilon$ and various $\ell$ values. The data in Table 3 clearly demonstrate that the center-of-mass energy $E_{\text{CM}}$ increases gradually as the black hole approaches the maximally spinning limit ($\epsilon\to 0$). For a fixed value of $\epsilon$, $E_{\text{CM}}$ rises with increasing values of the LQG parameter $\ell$. This behavior contrasts with the Kerr-Newman case Wei:2010vca , where $E_{\text{CM}}$ decreases as the charge increases, but aligns with the Kerr-Sen black hole scenario Wei:2010gq , in which $E_{\text{CM}}$ increases with increasing charge. Additionally, when $\ell=0$, the spacetime corresponds to a Kerr black hole, and the results in Table 3 are consistent with those reported in Ref. Jacobson:2009zg .
$\ell$ $\epsilon=0.1$ $\epsilon=0.05$ $\epsilon=0.01$ $\epsilon=0.001$ $\epsilon=0.0001$
00 6.90064 8.24403 12.5353 22.6342 40.4856
$0.1$ 6.90387 8.24916 12.5464 22.6584 40.5315
$0.2$ 6.91397 8.26531 12.5816 22.7356 40.6787
$0.3$ 6.93246 8.29503 12.6475 22.8826 40.9597
$0.4$ 6.96237 8.34370 12.7590 23.1386 41.4553
$0.5$ 7.00982 8.42251 12.9495 23.5977 42.3625
$0.6$ 7.08872 8.55867 13.3125 24.5610 44.3473
$0.7$ 7.24973 8.86388 14.3815 28.7519 55.5305
Table 3: The center-of-mass energy per unit rest mass $E_{\text{CM}}/m$ with spin $a=a_{\text{e}}-\epsilon$, and, $L_{1}=L_{\text{max}},L_{2}=L_{\text{min}}$.
Lastly, we briefly assess the tolerance level of the critical angular momentum $L_{\text{c}}$ in the BSW mechanism. For simplicity, we assume $L_{1}=L_{\text{c}}-\delta L$ and $L_{2}=0$, which simplifies Eq. (38) to
$\frac{E^{2}_{\mathrm{CM}}}{2m^{2}}\approx 9.60695\times 10^{17}\delta L+\frac{3.82061\times 10^{18}}{\delta L}+1.68323\times 10^{17}.$
[…]
$\delta L\approx 3.82061\times 10^{-22}$
which represents the tolerance for the critical angular momentum $L_{\text{c}}$.